# Optimization procedure of the baffle of the GroundBIRD Telescope to mitigate stray light

### v1

@@ III-B The baffle aperture angle size dependence of the thermal loading from the baffle
In this section, the baffle aperture angle dependence of the thermal loading from the absorbers put on the baffle is examined. The inner wall of the baffle is entirely covered by radio-absorptive material (ECCOSORB AN-72) [28] bonded to the inner surfaces using Stycast 2850FTJ adhesive (Henkel Corporation) [29]. The absorber emits blackbody radiation corresponding to the ambient temperature of approximately 293 K. Therefore, the thermal emission from the baffle can be one of the external thermal loads. Since the location of the baffle is well within the Fraunhofer distance of $R=2D^{2}/\lambda$, that is about 45 m for 145 GHz and 70 m for 220 GHz, the near-field beam pattern must be evaluated to estimate the thermal loading from the baffle.
@@
where $\tau_{\nu}$ is the optical depth of the atmosphere at frequency $\nu$ and $\eta_{\mathrm{ap}}$ is the aperture efficiency calculated by GRASP. Since most of the atmospheric emission originates from the far field and its intensity distribution is nearly uniform across the field of view of each detector pixel, we assume that the effective area times the field of view equals $\lambda^{2}$. The thermal loading from the baffle and the atmospheric emission are compared in table I for the detector pixels labeled in Fig. 2. We assumed that the GroundBIRD is pointing 20 degrees from the zenith and atmospheric PWV is 1 mm which is the optimal value representing good weather conditions at the Teide observatory. It shows if we take $\Theta=15^{\circ}$, the thermal loading of the baffle is comparable to the atmospheric contribution. However, as far as $\Theta$ is larger than 18 degrees, the thermal loadings due to the baffle are less than 15 % of the atmospheric contribution. Fig. 4c shows the cross section of the near field beam pattern for pixel ID 5 crossing its peak power position and $\theta=0$ as an example of the pixels located at the edge of the focal plane. It shows that a significant portion of the main beam of this pixel overlaps with the baffle when $\Theta=15^{\circ}$ is selected. Since the power of the main beam starts to decrease drastically at $\theta\sim-15^{\circ}$, the thermal loading from the baffle also decreases drastically when the aperture angle of the baffle becomes wider than $\Theta>15^{\circ}$. It shows that at $\theta=-20^{\circ}$, the power of the near field main beam is as small as 20-20 dB from its peak power.
III-C The baffle aperture angle dependence of the far field beam pattern
@@
Fig. 4c shows that the near field main beam is interfered with a baffle significantly when $\Theta=15^{\circ}$. It indicates that the far field main beam pattern could be significantly disturbed when the aperture angle of the baffle is set to be as narrow as $\Theta=15^{\circ}$. To quantify how the quality and the full width at half maximum (FWHM) of the main beam is degraded by narrowing the aperture angle of the baffle, $\Theta$ dependence of the ellipticity of the simulated main beam was evaluated. A 2D Gaussian function fitted the main beam patterns:
@@
where $x$ and $y$ coordinates are set to azimuth and elevation directions, respectively, $(x_{c},y_{c})$ represents the position of maximum beam intensity, and $\delta$ is the angle of the major axis relative to $x$ coordinate. The primed coordinates are defined as $x^{\prime}=x-x_{c}$ and $y^{\prime}=y-y_{c}$. The ellipticities of the main beams are defined by $\sigma_{2}/\sigma_{1}$. The main beam shapes were maintained with a constant FWHM and ellipticity for the central pixel, regardless of the angles of the baffle aperture. Other than the central pixel, the main beam shapes deviate from circular even when the wide baffle is installed. The deviation from the circular is worst for Pixel ID 1 and 2. They have ellipticity of 95 %. When the baffle with $\Theta=18^{\circ}$ is installed, pixel ID 4 shows the worse ellipticity of 92 %. When the baffle with $\Theta=15^{\circ}$ is installed, pixel ID 2 shows the worse ellipticity of 69 %. We have checked that the FWHM of main beams do not change significantly for all analyzed pixels as far as $\Theta\geq 18^{\circ}$.
@@ IV Optimized Baffle Design
Based on the results presented in the previous section, the optimized baffle is designed with an aperture angle of approximately $18^{\circ}$ to achieve maximum performance. Fig. 6 shows a photograph of the optimized baffle.
Figure 6: A photograph of the optimized baffle.
Our design adopts a multi-stage cylindrical structure rather than the conical geometry adopted in the wide baffle, as shown in Fig. 6. A multi-stage is selected to facilitate the manual installation of the baffle. A cylindrical structure is selected to facilitate the installation of absorbers on the baffle wall (see Section III-B). Since the thermal loading from the baffle wall only depends on the area of the baffle wall perpendicular to the near-field beam, this structural change will not increase the thermal loading compared to a conical structure baffle with the same aperture diameter.
The optimized baffle geometry features are as follows. The baffle length, $l$, was set to be $l=82$ cm so as not to interfere with the dome. The diameter of the baffle was 54 cm. After lining with the absorber with thickness of 0.6 cm, the diameter of the baffle aperture, $D$, is reduced to $\sim$52.8 cm. Therefore, the aperture angle is $\Theta=\arctan\left(\frac{D/2}{l}\right)=\arctan\left(\frac{52.8/2}{82}\right)\sim 17.8^{\circ}$.
@@ V Performance Evaluation
V-A The stray light reductions
@@
We confirmed that the ghost images of the Moon are suppressed for the other 4 pixels mounted on the same chip as pixel ID 9, as expected from the simulation.
@@ V-B Performance of the Thermal loading
We evaluate the optical loading budget by comparing the NETs of two different baffle configurations (optimized and wide baffles). The methods of extracting NET using the Moon observations are given in our previous study[32]. The error bars for NET are derived from the root mean square of the white noise component in the power spectral density (PSD) over the frequency range of 100 – 500 Hz. The error bars for PWV are derived from PWV measurements at Teide Observatory [33, 34]. Fig. 8 shows the NET for various PWVs. There is a PWV dependence of NETs measured with the wide baffle. It confirms that the dominant noise source is the atmospheric emission in the case of the wide baffle. Since the NET obtained with optimized baffle is consistent with those obtained with the wide baffle, we conclude that the thermal loading from the optimization of the baffle design is negligible compared with the atmospheric contribution.
@@
This is confirmed for other active pixels mounted on the same chip as ID 9.
@@
We have presented the optimization procedures of the baffle mounted on the GroundBIRD telescope, which employs dual-mirror reflector telescopes installed in the cryostat. The key strategies were to reduce stray light contamination as much as possible while maintaining the quality of the main beam and keeping the thermal loading from the baffle significantly below the atmospheric thermal loading. Using quasi-optics simulations, we optimized the baffle’s aperture angle to suppress stray light without degrading the main beam quality. We confirmed through the Moon observations that the optimized baffle design works to eliminate stray light contamination as expected. Additionally, there was no degradation in NET, indicating minimal thermal impact. Although simulation and observational data suggest that the optimized baffle does not affect the main beam, further tests with a far-field point source would be beneficial. However, due to the absence of a bright celestial point source for GroundBIRD, alternative methods such as drone-mounted artificial sources may be necessary[30, 31]. This work not only improves the current performance of the GroundBIRD telescope but also offers practical guidance for future CMB experiments.
acknowledgments
This work was supported by the GroundBIRD collaboration. This work was supported by MEXT KAKENHI Grant Number JP18H05539 and JSPS KAKENHI Grant Numbers JP15H05743, JP20K20927, JP20KK0065, JP21H04485, JP21K03585, JP22H04913, and JP24H00224, JSPS Bilateral Program Numbers JPJSBP120219943 and JPJSBP120239919, and also supported by JSPS Core-to-Core Program JPJSCCA20200003. M.T. also acknowledges support from Graduate Program on Physics for the Universe (GP-PU), Tohoku University, and JST SPRING, Grant Number JPMJSP2114. We thank Victor González Escalera, Ángeles Pérez de Taoro (the Instituto de Astrofísica de Canarias), and the staff of Teide Observatory for supporting the maintenance and operation of GroundBIRD. M.T. gratefully acknowledges the Machine shop, School of Science, Tohoku University and G. Nakamura for their valuable consultation and technical assistance in the baffle fabrication process.

### v2

@@ III-B Baffle Aperture Angle Size Dependence Of the Thermal Loading From the Baffle
In this section, the baffle aperture angle dependence of the thermal loading from the absorbers put on the baffle was examined. The inner wall of the baffle is entirely covered by radio-absorptive material (ECCOSORB AN-72) [33] bonded to the inner surfaces using Stycast 2850FTJ adhesive (Henkel Corporation) [34] (see Fig. 5b). The absorber emits blackbody radiation corresponding to the ambient temperature of approximately 293 K. Therefore, the thermal emission from the baffle can be one of the external thermal loads. Since the location of the baffle is well within the Fraunhofer distance of $R=2D^{2}/\lambda$, that is about 45 m for 145 GHz and 70 m for 220 GHz, the near-field beam pattern must be evaluated to estimate the thermal loading from the baffle.
@@
where $\tau_{\nu}$ is the optical depth of the atmosphere at frequency $\nu$ and $\eta_{\mathrm{ap}}$ is the aperture efficiency calculated by GRASP. Since most of the atmospheric emission originates from the far field and its intensity distribution is nearly uniform across the field of view of each detector pixel, we assume that the effective area times the field of view equals $\lambda^{2}$. The thermal loading from the baffle and the atmospheric emission are compared in Table I for the detector pixels labeled in Fig. 3. We assumed that GroundBIRD was pointing 20 degrees from the zenith and atmospheric PWV is 1 mm which is the optimal value representing good weather conditions at the Teide observatory. It shows that if we take $\Theta=15^{\circ}$, the thermal loading of the baffle is comparable to the atmospheric contribution. However, as far as $\Theta$ is larger than 18 degrees, the thermal loading due to the baffle is less than 15 % of the atmospheric contribution. Fig. 7 shows the cross section of the near field beam pattern for pixel ID 5 crossing its peak power position and $\theta=0$ as an example of the pixels located at the edge of the focal plane. It shows that a significant portion of the main beam of this pixel overlaps with the baffle when $\Theta=15^{\circ}$ is selected. Since the power of the main beam starts to decrease drastically at $\theta\sim-15^{\circ}$, the thermal loading from the baffle also decreases drastically when the aperture angle of the baffle becomes wider than $\Theta>15^{\circ}$. It shows that at $\theta=-20^{\circ}$, the power of the near field main beam is as small as 20-20 dB relative to its peak power.
III-C Baffle Aperture Angle Dependence of the Far Field Beam Pattern
@@
Figure 9: Two-dimensional beam patterns of pixel ID 5 at 145 GHz for different baffle aperture angles. The coordinates $(x^{\prime},y^{\prime})$ represent the angular offset from the beam center in azimuth and elevation directions. The red contours show the ellipse of the fitted 2D Gaussian function defined by $(x^{\prime}\cos\delta+y^{\prime}\sin\delta)^{2}/\sigma_{1}^{2}+(-x^{\prime}\sin\delta+y^{\prime}\cos\delta)^{2}/\sigma_{2}^{2}=1$, which represents the beam shape. (a) $\Theta=15^{\circ}$ (ellipticity 0.13). (b) $\Theta=18^{\circ}$ (ellipticity 0.03). (c) Wide baffle $\Theta=27^{\circ}$ (ellipticity 0.02).
@@
Fig. 7 shows that the near field main beam is diffracted by the baffle significantly when $\Theta=15^{\circ}$. It indicates that the far field main beam pattern could be significantly disturbed when the aperture angle of the baffle is set to be as narrow as $\Theta=15^{\circ}$. To quantify how the quality and the full width at half maximum (FWHM) of the main beam is degraded by narrowing the aperture angle of the baffle, $\Theta$ dependence of the ellipticity of the simulated main beam was evaluated. The main beam patterns were fitted with a 2D Gaussian function:
@@
where $x$ and $y$ coordinates are set to azimuth and elevation directions, respectively, $(x_{c},y_{c})$ represents the position of maximum beam intensity, and $\delta$ is the angle of the major axis relative to $x$ coordinate. The primed coordinates are defined as $x^{\prime}=x-x_{c}$ and $y^{\prime}=y-y_{c}$. The ellipticities of the main beams are defined by $|\sigma_{1}-\sigma_{2}|/(\sigma_{1}+\sigma_{2})$, which ranges from 0 (perfectly circular) to 1 ($\sigma_{1}>>\sigma_{2}$, or vice versa). Fig. 9 shows representative examples of the two-dimensional beam patterns for pixel ID 5 under three different baffle configurations. The fitted ellipses clearly visualize the degradation of beam circularity as the baffle aperture narrows from $27^{\circ}$ to $15^{\circ}$, with the ellipticity increasing from 0.02 to 0.13. Table II summarizes the ellipticity values for selected detector pixels at aperture angles of $15^{\circ}$, $18^{\circ}$, and $27^{\circ}$ (the wide baffle) to demonstrate the trend across different aperture configurations. The main beam shapes were maintained with a constant FWHM and ellipticity for the central pixel (pixel ID 0), regardless of the angles of the baffle aperture. The main beam shapes deviate from a perfect circle for all pixels even when the wide baffle is installed. With the wide baffle, the deviation from a circle is the worst for Pixel ID 1, which has an ellipticity of 0.03. When the baffle with $\Theta=18^{\circ}$ is installed, pixel IDs 1 and 4 show the worst ellipticity of 0.04. When the baffle with $\Theta=15^{\circ}$ is installed, pixel ID 2 shows the worst ellipticity of 0.18. We have checked that the FWHM of main beams does not change significantly for all analyzed pixels as far as $\Theta\geq 18^{\circ}$.
@@
TABLE II: Ellipticity of main beams for different baffle aperture angles.
Pixel ID $\Theta=15^{\circ}$ $\Theta=18^{\circ}$ The wide baffle
0 0.01 0.01 0.02
1 0.10 0.04 0.03
2 0.18 0.01 0.02
3 0.07 0.01 0.02
4 0.11 0.04 0.02
5 0.13 0.03 0.02
6 0.15 0.03 0.02
7 0.15 0.03 0.02
8 0.13 0.03 0.02
9 0.02 0.01 0.01
@@ IV Optimized Baffle Design
Based on the results presented in the previous section, the optimized baffle was designed with an aperture angle of approximately $18^{\circ}$ to achieve maximum performance. Fig. 10 shows a photograph of the optimized baffle.
Figure 10: A photograph of the optimized baffle. (a) External view showing the multi-stage cylindrical structure. (b) Internal view showing the radio-absorptive material (ECCOSORB AN-72) covering the inner wall. The absorber is bonded to the aluminum surfaces using Stycast 2850FTJ adhesive.
Our design adopted a multi-stage cylindrical structure rather than the conical geometry adopted in the wide baffle, as shown in Fig. 10a. The baffle is constructed of multiple pieces connected by flanges to facilitate installation. A cylindrical structure was selected to facilitate the installation of absorbers on the baffle wall (see Section III-B). The internal view of the baffle is shown in Fig. 10b. Since the thermal loading from the baffle wall only depends on the area of the baffle wall perpendicular to the near-field beam, this structural change does not increase the thermal loading compared to a conical structure baffle with the same aperture diameter.
The optimized baffle geometry features are as follows. The baffle length, $l$, was set to be $l=82$ cm so as not to interfere with the dome. The diameter at the top of the baffle was 54 cm. After lining with the absorber with a thickness of 0.6 cm, the diameter at the top of the baffle aperture, $D$,was reduced to $\sim$52.8 cm. Therefore, the aperture angle is $\Theta=\arctan\left(\frac{D/2}{l}\right)=\arctan\left(\frac{52.8/2}{82}\right)\sim 17.8^{\circ}$.
@@ V Performance Evaluation
V-A Stray Light Reductions
@@
We confirmed that the ghost images of the Moon were suppressed for the other 4 pixels mounted on the same chip as pixel ID 9, as expected from the simulation.
@@ V-B Performance of the Thermal Loading
We evaluated the optical loading budget by comparing the NETs of the two different baffle configurations (optimized and wide baffles). The methods of extracting NET using Moon observations are given in our previous study[35]. The error bars for NET are derived from the root mean square of the white noise component in the power spectral density (PSD) over the frequency range of 100 – 500 Hz. The error bars for PWV are derived from PWV measurements at Teide Observatory [36, 37]. Fig. 12 shows the NET for various PWVs. There is a PWV dependence of NETs measured with the wide baffle. It confirms that the dominant noise source is the atmospheric emission in the case of the wide baffle. The NET obtained with optimized baffle is consistent with that obtained with the wide baffle when $\mathrm{PWV}=1.9$ mm. This confirms that the noise loading due to thermal emission from the absorbers attached to the baffle is negligibly small compared with the noise due to atmospheric emission at $\mathrm{PWV}=1.9$ mm. According to the GroundBIRD forecast paper[38], the scientific goal of the GroundBIRD is achieved by three-year observations under atmospheric conditions of $\mathrm{PWV}=4$ mm. The average PWV at the Teide Observatory from 2023 July to 2024 August was 4.3 mm[39]. We concluded that the thermal loading due to the optimization of the baffle design has a negligible impact on achieving the scientific goal of the GroundBIRD.
@@
This was confirmed for other active pixels mounted on the same chip as ID 9.
@@ VI Discussion and Conclusion
We have presented the optimization procedures of the baffle mounted on the GroundBIRD telescope, which employs a dual-mirror reflector telescope installed in the cryostat. The key strategies were to reduce stray light contamination as much as possible while maintaining the quality of the main beam and keeping the thermal loading from the baffle significantly below the atmospheric thermal loading. Using quasi-optics simulations, we optimized the baffle’s aperture angle to suppress stray light without degrading the main beam quality. We confirmed through Moon observations that the optimized baffle design works to eliminate stray light contamination as expected. Additionally, there was no degradation in NET, indicating minimal thermal impact. Although simulation and observational data suggest that the optimized baffle does not affect the main beam, further tests with a far-field point source would be beneficial. However, due to the absence of a bright celestial point source for GroundBIRD, alternative methods such as drone-mounted artificial sources may be necessary[40, 41]. This work not only improves the current performance of the GroundBIRD telescope but also offers practical guidance for future CMB experiments.
Acknowledgment
This work was supported by the GroundBIRD collaboration. This work was supported by MEXT KAKENHI Grant Number JP18H05539 and JSPS KAKENHI Grant Numbers JP15H05743, JP20K20927, JP20KK0065, JP21H04485, JP21K03585, JP22H04913, and JP24H00224, JSPS Bilateral Program Numbers JPJSBP120219943 and JPJSBP120239919, and also supported by JSPS Core-to-Core Program JPJSCCA20200003. M.T. also acknowledges support from Graduate Program on Physics for the Universe (GP-PU), Tohoku University, and JST SPRING, Grant Number JPMJSP2114. We thank Victor González Escalera, Ángeles Pérez de Taoro (the Instituto de Astrofísica de Canarias), and the staff of Teide Observatory for supporting the maintenance and operation of GroundBIRD. M.T. gratefully acknowledges the Machine shop, School of Science, Tohoku University and G. Nakamura for their valuable consultation and technical assistance in the baffle fabrication process. We thank anonymous referees for providing constructive comments. These were very helpful in revising our paper.